# Hierarchy Pruning for Unseen Domain Discovery in Predictive Healthcare

### v2

@@ Abstract
Healthcare providers often divide patient populations into cohorts based on shared clinical factors, such as medical history, to deliver personalized healthcare services. This idea has also been adopted in clinical prediction models, where it presents a vital challenge: capturing both global and cohort-specific patterns while enabling model generalization to unseen domains. Addressing this challenge falls under the scope of domain generalization (DG). However, conventional DG approaches often struggle in clinical settings due to the absence of explicit domain labels and the inherent gap in medical knowledge. To address this, we propose UdonCare, a hierarchy-guided method that iteratively divides patients into latent domains and decomposes domain-invariant (label) information from patient data. Our method identifies patient domains by pruning medical ontologies (e.g. ICD-9-CM hierarchy). On two public datasets, MIMIC-III and MIMIC-IV, UdonCare shows superiority over eight baselines across four clinical prediction tasks with substantial domain gaps, highlighting the untapped potential of medical knowledge in guiding clinical domain generalization problems.
@@ 1 Introduction
The digitization of clinical data, notably electronic health records (EHR), has transformed healthcare by enabling efficient computational analysis. Current deep learning techniques have also achieved significant gains in diagnosis, mortality, and readmission prediction tasks (Poulain & Beheshti, 2024; Jiang et al., 2024). Still, these models trained on the training (source) data often suffer performance drops when applied to the test (target) data under domain shifts, that is, distributional changes across patient groups, such as data from different hospitals (Perone et al., 2019; Koh et al., 2021). Consequently, handling domain shifts is a prerequisite for alleviating performance degradation in clinical predictive models (Yang et al., 2023a; Wu et al., 2023). It also aligns with the objective of most domain generalization (DG) methods, such as meta-learning (Balaji et al., 2018; Dou et al., 2019), adversarial learning (Ganin et al., 2016; Li et al., 2018b), and latent-domain techniques (Matsuura & Harada, 2020; Wu et al., 2023).
@@
In this work, we focus on tackling DG problem in clinical settings, whereas most recent models have been developed for image classification. However, directly transferring these regular DG methods will encounter two clinical-specific obstacles: (1) Domain IDs, which are naturally defined in image datasets (e.g., dog & cat), are unseen in most EHR datasets, but most DG solutions require the presence of domain IDs (Wu et al., 2023). Some studies treat each patient as unique domain (Dou et al., 2019; Yang et al., 2023a), which is overly fine-grained and unstable. Others rely on broader categorizations (e.g. institute & admission period), which overlook clinical heterogeneity (Zhang et al., 2021a; Guo et al., 2022). (2) Even though some DG methods do not rely on domain IDs (Arjovsky et al., 2019; Liu et al., 2021b), they overlook clinical semantics. For instance, Matsuura & Harada (2020); Wu et al. (2023) cluster patient features to form latent domains, but the resulting partitions are highly sensitive to training data. In practice, patient groups can vary significantly, even over longer admission periods, since they reflect only feature-level similarity without capturing the progression of medical concepts. Hence, it is crucial to construct robust domains with explicit definitions grounded in clinical relevance.
@@
To address these challenges, we explore the following research question: Instead of assuming the presence of domain IDs, can we leverage medical knowledge to guide models in discovering domains that are both adaptive and clinically meaningful? In most hospitals, visiting patients are treated based on their medical history, which is expressed through medical concepts shown in their admissions. For instance, when dealing with heart failure patients, hospitals may categorize heart failure as a distinct domain or group it with other cardiovascular diseases. Similarly, in medical ontologies like ICD-9-CM, heart failure corresponds to a leaf node under a higher-level node grouping cardiovascular disease, and such hierarchical relation motivates us to use a pruning algorithm to identify appropriate ancestor nodes for domain partitioning. Building upon this perspective, our work focuses on knowledge-guided partitioning of medical concepts, ensuring alignment with clinical semantics while maintaining flexibility for adaptive generalization.
To this end, we propose UdonCare, a framework that integrates medical ontologies into an iterative domain discovery process, enabling domain distinction at varying levels of abstraction. It ensures the discovered domains remain consistent with clinical reasoning while supporting robustness against distribution shifts. Specifically, a pruning algorithm is developed to merges similar concepts on hierarchies and generate soft labels as domain IDs for patients. To explicitly remove domain features from patients, we leverage a mutual learning network, which learns domain-invariant (label) representations upon the orthogonal factorization. Finally, domain assignments and feature extraction are updated jointly through an iterative collaborative inference mechanism, allowing the pruning module to adapt domain categorization according to input data and task settings. Our main contributions are enumerated as follows:
@@
  • To the best of our knowledge, this is the first work using medical ontologies to tackle clinical DG problems. It reveals the potential of medical ontologies in finding latent domains for handling covariates, rather than serving as feature enrichment (Lu et al., 2021; Jiang et al., 2024).
@@
  • UdonCare shows accurate prediction across four vital predictive tasks on two public datasets, outperforming both clinical DG baselines (Yao et al., 2022; Wu et al., 2023) and regular DG baselines. UdonCare boosts the AUPRC score by $5-20\%$ over the best baselines.
@@
  • We conduct detailed analyses to show that UdonCare addresses domain shifts through accurate domain partitioning and invariant feature learning, without sacrificing computational overhead.
@@ 2 Preliminary
EHR Dataset. Given an EHR dataset $\mathcal{S}$, the $i$-th patient’s data $\mathbf{x}^{(i)}$ consists of a longitudinal sequence of visits $\{V^{(i)}_{1},V^{(i)}_{2},\dots,V^{(i)}_{T}\}$. We omit the patient index $i$ to illustrate our method using single patient data $\mathbf{x}$. Medical codes $c_{i}$ in admissions can be also categorized into $K$ distinct feature keys. In this work, we identify feature keys from the vocabulary of medical concepts $\mathcal{C}\in\{\mathcal{D},\mathcal{P},\mathcal{M}\}$, where $\mathcal{D},\mathcal{P},\mathcal{M}$ denote the sets of diseases, procedures, and drugs, respectively.
EHR Predictive Models. For clinical prediction, models trained on EHR data typically aim to predict clinical outcomes $\mathbf{y}\in\{0,1\}^{d}$ at a future visit $V_{t+1}$, where $d$ is the number of labels. To learn temporal changes of feature key $k$, most studies develop a feature extractor $f_{\phi,k}(\cdot):\mathbf{x}_{k}\mapsto\mathbf{p}_{k}$ to encode the admission sequence $\mathbf{x}_{k}$ into a patient-level embedding $\mathbf{p}_{k}$, which can then be concatenated with embeddings from other keys or used directly for downstream predictions:
$\mathbf{p}=\mathbf{p}_{1}\oplus\dots\oplus\mathbf{p}_{k}\oplus\dots\oplus\mathbf{p}_{K},\;\text{where}\;\mathbf{p}_{k}=f_{\phi,k}(\mathbf{x}_{k})\in\mathbb{R}^{h}.$ (1)
Here, $f_{\phi,k}$ is the encoder for feature key $k$, $\oplus$ denotes vector concatenation across $K$ feature keys, and $h$ is the embedding dimension. Most studies assume that these learned embeddings effectively capture patient’s medical history, and $\mathbf{p}$ is then passed to a label predictor (e.g. MLP).
Concept-Specific Hierarchy. In EHR data, certain medical concept $c_{i}\in\mathcal{C}$ always originates from a hierarchical encoding system, such as ICD-9 (Organization et al., 1988) and ATC (Nahler & Nahler, 2009) codes. We define a concept-specific hierarchy $\mathcal{H}$ of $H$ levels, and denote $n^{(h)}_{i}$ as the $i$-th node on level $h$. Leaf nodes at level $H$ represent actual codes via the mapping $m:c_{i}\mapsto n^{(H)}_{i}$ with node feature $\mathbf{e}_{i}$ stored in $f_{\phi,k}$, and $\mathrm{Desc}(\cdot)$ denotes the set of descendant nodes. Note that, the root node $n^{(1)}_{1}$ at the top level subsumes all nodes in $\mathcal{H}$, and any two leaf nodes $n^{(H)}_{i}$ and $n^{(H)}_{j}$ share at least one common ancestor. In this work, we only focus on the disease hierarchy (i.e. ICD-9-CM), as incorporating treatments and drugs yields only marginal gains (see Appendix F). However, UdonCare can be extended to incorporate additional feature keys or ontologies if needed.
Figure 1: The Overall Framework of UdonCare. The forward structure adds a domain pathway for mutual learning, extending beyond the backbone pathway of conventional predictive models. During training, we first feed patient data $\mathbf{x}$ into the backbone pathway, which learns patient features $\mathbf{p}$ through $f_{\phi,k}(\cdot)$ and produces the output prediction $\hat{y}_{p}$. In parallel, we obtain $\hat{y}_{h}$ from invariant features $\mathbf{h}$ along the domain pathway by applying $\mathrm{DiscoveryAlgo}(\cdot)$, $g_{\theta,k}(\cdot)$, and $h(\cdot)$. Here we iteratively adapt latent domains in $\mathbf{M}$ and update parameters on both pathways by ground truths $y$.
@@ 3 Methodology
To generalize $f_{\phi}(\cdot)$ on target data despite domain shifts, we develop a hierarchy-guided framework that iteratively divides patients into latent domains and decompose domain-invariant features for downstream health risk predictions. It iteratively operates two main steps:
Step 1: Develop a pruning algorithm for medical hierarchies to discover latent domains;
Step 2: Learn invariant (label) information by factorizing patient features in projection space.
Figure 1 presents a workflow of UdonCare. In general, we aim to show how medical knowledge can guide domain generalization for clinical prediction, rather than merely augment features.
3.1 Step 1: Hierarchy-Guided Domain Discovery
Figure 2: A simple illustration of hierarchy-guided domain discovery.
While domain IDs are unobserved in EHR datasets, it is intuitive that patients with similar medical histories (concepts) often belong to the same domain. We can divide patient cohorts by treating the multi-hot vector $\mathbf{v}_{t}$ from admissions as a soft domain label. However, the number of latent domains grows exponentially with the larger vocabulary size $|\mathcal{C}|$ (i.e. $2^{|\mathcal{C}|}$). Here, we design a hierarchy-guided domain discovery algorithm that assigns and updates domain IDs for patients’ admissions. Our goal is to prune overly fine-grained nodes, thereby forming a smaller set of ancestors that still covers all concepts. Given a set of patient samples $\{\mathbf{x}^{(i)}\}_{i=1}^{N_{\mathrm{tr}}}$, we construct a assignment matrix $\mathbf{M}$ to query domain IDs via
@@ 3.1 Step 1: Hierarchy-Guided Domain Discovery
where $N_{\mathrm{tr}}$ is the number of training patients, $|\mathcal{C^{{}^{\prime}}}|$ is the pruned vocabulary size of medical concepts. It merges fine-grained codes into fewer, higher-level clusters, allowing patients with or without a particular disease to occupy different latent domains as needed.
Initialization on Domain IDs. Following the previous settings (Lu et al., 2021; Jiang et al., 2024), we only focus on patients with multiple admission, where there are $T\geq 2$ records. This sequence should be then converted into a single vector that consolidates all prior visits. Hence, we aggregate patient data $\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}$ into a unified representation $\mathbf{X}=\bigvee_{t=1}^{T}\mathbf{x}_{t}$, where medical concepts shown in each admission are merged to form a comprehensive medical history. Note that, $\mathbf{X}$ is the most fine-grained domain assignment, which initializes $\mathrm{DiscoveryAlgo}(\cdot)$.
Initialization on Node Features. After getting $\mathbf{X}$ for each patient, concept-specific ontology $\mathcal{H}$ with node features are required to decide whether fine-grained medical concepts group diseases into higher-level clusters or are preserved. We initialize leaf-node features $\mathbf{e}_{i}\in\mathbb{R}^{h}$ by: (1) for present code $c_{i}$ in $\mathcal{S}$, $\mathbf{e}_{i}$ is initialized from embedding layer $\text{E}(e_{1},\dots,e_{|\mathcal{C}|})$ in $f_{\phi,k}(\cdot)$; (2) for absent code $c_{i}$ in $\mathcal{S}$, $\mathbf{e}_{i}$ is its embedding of entity name through ClinicalBERT (Huang et al., 2019). The feature of a parent node $\mathbf{e}^{(h-1)}_{n_{i}}$ is then computed from the embeddings of its descendants at level $h$:
$\mathbf{e}^{(h-1)}_{i}\;:=\;\frac{1}{|\mathrm{Desc}(n^{(h-1)}_{i})|}\sum_{n\in\mathrm{Desc}(n_{i})}\mathbf{e}^{(h)}_{n}$ (3)
It extends $\text{E}(e_{1},\dots,e_{|\mathcal{C}|})$ to $\text{E}(e_{1},\dots,e_{|\mathcal{H}|})$ over the entire hierarchy. Still, it fails to capture the hierarchical distances in $\mathcal{H}$. For example, two codes might be totally different despite sharing the same parent node. Therefore, we mimic the principle of hierarchical clustering (Johnson, 1967) by propagating node features upward based on feature similarity. For each most similar node pair $(d_{i},d_{j})$, their lowest common ancestor $\mathrm{LCA}(d_{i},d_{j})$ is updated by averaging their embeddings:
@@
This process continues until the maximum similarity among remaining pairs falls below a threshold $\rho=0.3$, yielding hierarchy-aware embeddings that integrate structural positions with features. See Appendix H for more details about the information flow.
Node Scoring. For each node $n\in\mathcal{H}$, we define $S(n)$ to identify which node is a good “candidate” for final selection. Motivated by the idea of information gains (Song & Ying, 2015), $S(n)$ involves three indicators, coverage $\mathrm{cov}(n)$, purity $\mathrm{pur}(n)$, and depth $\mathrm{dep}(n)$, via
$S(n)$ $=\alpha\cdot\exp(\mathrm{pur}(n))+(1-\alpha)\bigl(\mathrm{cov}(n)\times\mathrm{dep}(n)\bigr)$
$=\alpha\cdot\exp\Bigl(\mathbb{E}_{m\in\mathcal{M}}[\mathrm{sim}(\mathbf{e}_{n},\mathbf{e}_{m})]\Bigr)+(1-\alpha)(\frac{|\mathcal{M}|}{|\mathcal{L}|}\cdot\frac{h}{H})$ (5)
@@
  • If $S(p)>\max(\{S(c_{i})\}^{r})$, we include parent node $p$ and exclude its children.
  • If $S(p)<\min(\{S(c_{i})\}^{r})$, we discard parent node $p$ and select all children.
@@
  • Otherwise, we tentatively discard $p$ but mark it for further resolution in the next step.
After the first iteration on the score matrix, a candidate subset $\mathcal{C}_{0}$ will be generated. However, such a result can be considered as a local optimum solution, since marked candidates still require further evaluation. Here a list of tuples $A$ of length $N$ is adopted to trade off each flagged parent-child pair $A[n]:=(p,\{c_{1},\dots,c_{r}\})$ to find an optimal result via a chosen search strategy.
@@
Domain Searching. Given $N$ flagged pairs, the complexity of a typical search approach grows exponentially (i.e. $O(2^{N})$) in the number of pairs. To search near-optimal pruning results, we employ a rectified Beam-Search algorithm (Lowerre, 1976) for more efficient optimization. We compute a global clustering metric, Silhouette Score (Shahapure & Nicholas, 2020), to evaluate how well each pruned node separates its assigned leaves from others. For each flagged pairs, we either (1) unify them by including the parent or (2) retain the children as distinct pruned nodes, whichever achieves higher scores. This process iterates over all flagged nodes in order by updating pruning subsets in $\{\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{N}\}$ to make selection $\mathcal{C}^{\prime}$ with the optimal evaluation result.
Domain Decision. To this end, we obtain an updated vocabulary $\mathcal{C}^{\prime}\subseteq\mathcal{H}$ of selected higher-level nodes and update the domain-assignment matrix $\mathbf{M}$ by linking each patient’s admission records (originally from $\mathcal{C}$) to these pruned clusters in $\mathcal{C}^{\prime}$, where $|\mathcal{C}^{\prime}|\leq|\mathcal{C}|$. If patient $p^{(i)}$ has at least one leaf code $d$ that descends from the pruned node $p_{j}\in\mathcal{C}^{\prime}$, we update $\mathbf{M}[i,j]=1$; otherwise, $\mathbf{M}[i,j]=0$. Note that, the output matrix $\mathbf{M}$ define domain categorization in terms of node features.
3.2 Step 2. Mutual Forward Learning

### v3

# Discovering Hierarchy-Grounded Domains with Adaptive Granularity for Clinical Domain Generalization

@@ Abstract
Domain generalization has become a critical challenge in predictive healthcare, where different patient groups often exhibit shifting data distributions that degrade model performance. Still, regular domain generalization approaches often struggle in clinical settings due to (1) the absence of domain labels and (2) the lack of clinical insight integration. To address these challenges in healthcare, we aim to explore how medical ontologies can be used to discover dynamic yet hierarchy-grounded patient domains, a partitioning strategy that remains under-explored in prior work. Hence, we introduce UdonCare, a hierarchy-pruning method that iteratively divides patients into latent domains and retrieve domain-invariant (label) information from patient data. On two public datasets, UdonCare shows superiority over eight baselines across four representative clinical prediction tasks with substantial domain gaps, highlighting the potential of medical knowledge for enhancing model generalization.
Machine Learning, ICML
@@ 1 Introduction
The digitization of clinical data, notably electronic health records (EHR), has transformed healthcare by enabling efficient computational analysis. Current deep learning techniques have also achieved significant gains in diagnosis, mortality, and readmission prediction tasks (Poulain and Beheshti, 2024; Jiang et al., 2024). Still, these models trained on a set of training samples (source) often suffer performance drops when applied to a new test set (target) in a different domain. Domain shift denotes distributional changes between patient groups, such as data from different hospitals or time periods (Perone et al., 2019; Koh et al., 2021). Consequently, handling domain shift is a prerequisite for alleviating performance degradation in clinical predictive models (Yang et al., 2023a; Wu et al., 2023). It also aligns with the objective of most domain generalization (DG) methods, such as meta-learning (Balaji et al., 2018; Dou et al., 2019), feature alignment (Ganin et al., 2016; Li et al., 2018b), and latent-domain techniques (Matsuura and Harada, 2020; Wu et al., 2023).
@@
In this work, we study DG problems in healthcare applications, which poses unique challenges beyond conventional DG scenarios: (1) Domain IDs (labels) are usually unseen in EHR datasets, but most DG solutions require the presence of domain IDs (Wu et al., 2023). Thus, some studies treat each patient as a unique domain (Dou et al., 2019; Yang et al., 2023a), which is overly fine-grained and unstable for model training, and others rely on broader partitions (e.g. institute & admission year), which overlook clinical heterogeneity (Zhang et al., 2021a; Guo et al., 2022). (2) Even though some DG methods do not rely on domain IDs (Arjovsky et al., 2019; Liu et al., 2021b), they overlook clinical semantics. For instance, prior DG methods (Matsuura and Harada, 2020; Wu et al., 2023) cluster instance features to form latent domains, but the resulting partitions are highly sensitive to training data. In practice, patient groups vary significantly, especially over longer admission periods. Existing DG methods can capture feature-level similarity, but not the progression of medical concepts. Hence, it is crucial to construct robust domains with explicit definitions grounded in clinical relevance.
@@
To address these challenges, we explore the following research question: Instead of assuming the presence of domain IDs, can we leverage medical knowledge to guide models in discovering domains that are both adaptive and clinically meaningful? In hospitals, visiting patients are treated based on their medical history, which includes medical concepts such as diseases or medications. For instance, when dealing with heart failure patients, hospitals may categorize heart failure as a distinct domain or group it with other cardiovascular diseases. Similarly, heart failure corresponds to a leaf node under a higher-level node grouping cardiovascular diseases in ICD-9 ontologies (Organization and others, 1988), which motivates us to use a pruning algorithm to identify appropriate ancestor nodes for domain partitions.
However, translating medical ontologies into a form suitable for latent domain discovery is not a simple problem. Existing DG methods identify latent domains by clustering patient-level representations in vector spaces (Matsuura and Harada, 2020; Wu et al., 2023), which precludes the direct incorporation of ontology hierarchies. Meanwhile, medical ontologies are commonly used to enrich patient representations via graph neural networks (Jiang et al., 2024), but these ontologies are rarely explored as structural rules for discovering latent domains. Thus, we aim to bridge this disconnect and leverage ontologies not as auxiliary features, but as principled constraints that shape domain formation in a clinically meaningful way.
To this end, we introduce UdonCare, a framework that discovers Unseen domains with adaptive granularity to enhance model generalization in predictive healthcare. It ensures domain partitions remain consistent with clinical semantics while maintaining flexibility for adaptive generalization. Concretely, a pruning algorithm is developed to merges similar concepts on hierarchies and generate soft labels as domain IDs for patients. To explicitly remove covariates from patients, we propose a mutual learning network, which only retrieves label information upon the orthogonal factorization for final predictions. During training, model parameters are jointly optimized in an iterative manner across epochs, enabling the pruning module to adjust domain partitions based on the evolving latent embeddings. Our main contributions are enumerated as follows:
@@
  • To the best of our knowledge, this is the first work using medical ontologies to tackle clinical DG problems. It reveals the potential of medical ontologies for discovering latent domains for handling covariates, rather than serving as feature enrichment.
@@
  • UdonCare shows robustness across four predictive tasks on two public datasets, outperforming both clinical DG baselines (Yao et al., 2022; Wu et al., 2023) and regular DG baselines. For example, our model boosts the AUPRC score up to 11% over the best baselines.
@@
  • We extend more analyses to show that UdonCare handles domain shifts through hierarchy-grounded domain partitions and accurate invariant feature learning, without sacrificing computational overhead.
2 Related Work
Domain Generalization (DG).  DG is pursuing adjusted models which are specially designed to remove domain-covariate features from hidden representation (Muandet et al., 2013). Most studies are dedicated to solve performance drops on target domain across diverse scenarios like image classification (Zhou et al., 2022; Ding et al., 2022), and they can be categorized as three different ways: (i) An intuitive way is to minimize the empirical source risk, either domain alignment (Ganin et al., 2016; Li et al., 2018b; Zhao et al., 2020) and invariant learning technique (Liu et al., 2021a; Zhang et al., 2022a; Wang et al., 2022) aim to convey little domain characteristics to acquire task-specific features. (ii) Contrastive learning (Kim et al., 2021; Jeon et al., 2021; Yao et al., 2022) becomes an alternative for data augmentation, studies employed the contrastive loss function to reduce the gap of representation distribution in one category. (iii) Meta-learning (Balaji et al., 2018; Dou et al., 2019) and ensemble-learning (Cha et al., 2021; Chu et al., 2022) approaches handle domain shifts through dynamic loss functions. However, they predefine either numerous or few domains in clinical settings (Wu et al., 2023), which motivates us to design a precise and efficient way to discover latent domains from learnable parameters.
DG in Clinical Prediction.  Prior works have shown that EHR predictive models often suffer performance drops when transferred to unseen patients with different data distributions (Perone et al., 2019; Koh et al., 2021). Some studies (Zhao et al., 2020; Zhang et al., 2021b) adopt feature adaptation to align domain shifts across multiple hospitals and different time periods (Reps et al., 2022; Zhang et al., 2022b). Recently, more works consider model generalization by mitigating the patient-specific domain shifts (Guo et al., 2022; Hai et al., 2024), which offers a flexible alternative in more scenarios. Moreover, ManyDG (Yang et al., 2023a) learns invariant features by treating each patient as unique domain; SLDG (Wu et al., 2023) develops a mixture-of-domain method to divide patient domains by concept-level features. However, they address domain partition with simplifying heuristics such as linear dependencies (Li et al., 2020), while support from medical knowledge beyond training data remains unexplored in clinical DG problems.
Figure 1: The Overall Framework of UdonCare. The forward structure adds a domain pathway for mutual learning, extending beyond the backbone pathway of conventional predictive models. We feed patient data $\mathbf{x}$ into the backbone pathway, which learns patient features $\mathbf{p}$ through $f_{\phi,k}(\cdot)$ and produces the output prediction $\hat{y}_{p}$. In parallel, we obtain $\hat{y}_{h}$ along the domain pathway through $\mathrm{DiscoveryAlgo}(\cdot)$, $g_{\theta,k}(\cdot)$, and $h(\cdot)$. Here we can iteratively update domain partitions in $\mathbf{M}$ and update parameters on both pathways during training.
@@ 3 Preliminary
Clinical Prediction on EHR Data.   Each patient $\mathbf{x^{(i)}}$ consists of a longitudinal sequence of admissions $\{V^{(i)}_{1},V^{(i)}_{2},\dots,V^{(i)}_{T}\}$. We denote the entire set of medical concepts as $\mathcal{C}=\{c_{1},c_{2},...,c_{|\mathcal{C}|}\}$, and each admission contains a subset of $\mathcal{C}$. Most predictive models predict clinical outcomes $\mathbf{y}^{(i)}\in\{0,1\}^{d}$ at a future visit $V_{t+1}$, and they train a feature extractor $f_{\phi,k}(\cdot):\mathbf{x}_{k}\mapsto\mathbf{p}_{k}$ to encode $\mathbf{x}_{k}$ (omit $i$ in single patient) into patient embedding
$\mathbf{p}_{k}$ $=f_{\phi,k}(\mathbf{x}_{k})\in\mathbb{R}^{h},$ (1)
where $f_{\phi,k}$ is the encoder for feature key $k$, and $h$ denotes the dimension. When $f_{\phi}(\cdot)$ are trained on source training data sampled from distribution $P_{s}$, we often expect the learned model to also perform well on target data drawn from $P_{t}$ by assuming $P_{s}=P_{t}$. However, it is possible that these two distributions differs a lot $P_{s}\neq P_{t}$ due to spatial and temporal shifts, which motivates the need for DG solutions to factorize label information that remains invariant across clinical environments.
Concept-Specific Hierarchy.   In EHR data, certain medical concept $c_{i}\in\mathcal{C}$ always originates from a hierarchical encoding system, such as ICD (Organization and others, 1988) and ATC (Nahler and Nahler, 2009) codes. We define a concept-specific hierarchy $\mathcal{H}$ of $H$ levels, and denote $n^{(h)}_{i}$ as the $i$-th node on level $h$. Leaf nodes at level $H$ represent actual codes via the mapping $m:c_{i}\mapsto n^{(H)}_{i}$ with node feature $\mathbf{e}_{i}$ stored in $f_{\phi,k}$, and $\mathrm{Desc}(\cdot)$ denotes the set of descendant nodes. Note that, the root node $n^{(1)}_{1}$ at the top level subsumes all nodes in $\mathcal{H}$, and any two leaf nodes $n^{(H)}_{i}$ and $n^{(H)}_{j}$ share at least one common ancestor. In this paper, we focus on the disease hierarchy for method clarity, while more ontologies (treatments and medications) are explored in Appendix H.
@@ 4 Methodology
UdonCare iteratively operates two main steps (as shown in Figure 1): (1) developing a pruning algorithm for medical hierarchies to discover latent domains; (2) learning invariant features by factorizing patient embeddings in projection space. A notation table is also provided in Appendix C.
4.1 Hierarchy-Guided Domain Discovery
While domain IDs are unavailable in EHR data, it is intuitive that patients with similar medical diagnoses (concepts) often belong to the same domain. Naively partitioning patients by medical histories $\mathbf{x}$ implicitly induces soft domain labels, but leads to an exponential growth of latent domains with vocabulary size $|\mathcal{C}|$ (i.e., $2^{|\mathcal{C}|}$). To address this issue, we propose a hierarchy-guided domain discovery algorithm that prunes overly fine-grained nodes and assigns domain IDs at appropriate ancestor levels, forming a compact domain set that still covers all concepts (see Figure 2). An assignment matrix $\mathbf{M}$ can be generated to query domain IDs via
@@ 4.1 Hierarchy-Guided Domain Discovery
where $N_{\mathrm{tr}}$ is the number of training patients, $|\mathcal{C^{{}^{\prime}}}|$ is the pruned vocabulary size. It merges fine-grained codes into higher-level clusters, allowing patients with or without a particular disease to occupy different latent domains as needed.
Initialization on Domain IDs. Following the previous settings (Lu et al., 2021; Jiang et al., 2024), we only focus on patients with multiple admissions $T\geq 2$. This sequence should be then converted into a single vector that consolidates all prior visits. Hence, we aggregate patient data $\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}$ into a unified representation $\mathbf{X}=\bigvee_{t=1}^{T}\mathbf{x}_{t}$, where medical concepts shown in each admission are merged to form a accumulated medical history. Note that, $\mathbf{X}$ is the most fine-grained domain partition.
Initialization on Node Features. With $\mathbf{X}$ for each patient, concept-specific ontology $\mathcal{H}$ with node features are required to decide whether fine-grained medical concepts group diseases into higher-level clusters or are preserved. We initialize leaf-node features $\mathbf{e}_{i}\in\mathbb{R}^{h}$ by: (1) for present code $c_{i}$ in $\mathcal{S}$, $\mathbf{e}_{i}$ is initialized from embedding layer $\text{E}(e_{1},\dots,e_{|\mathcal{C}|})$ in $f_{\phi,k}(\cdot)$; (2) for absent code $c_{i}$ in $\mathcal{S}$, $\mathbf{e}_{i}$ is its embedding of entity name through ClinicalBERT (Huang et al., 2019). The feature of a parent node $\mathbf{e}^{(h-1)}_{n_{i}}$ is then computed from the embeddings of its descendants at level $h$:
$\mathbf{e}^{(h-1)}_{i}\;:=\;\frac{1}{|\mathrm{Desc}(n^{(h-1)}_{i})|}\sum_{n\in\mathrm{Desc}(n_{i})}\mathbf{e}^{(h)}_{n}$, (3)
which extends $\text{E}(e_{1},\dots,e_{|\mathcal{C}|})$ to $\text{E}(e_{1},\dots,e_{|\mathcal{H}|})$ over the entire hierarchy. Still, it fails to capture the hierarchical distances in $\mathcal{H}$. For example, two leaf nodes might have similar features without sharing the same parent node. Under this scenario, their lowest common ancestor (LCA) $n_{\text{LCA}}(d_{i},d_{j})$ could be an alternative reflecting such similarity. Hence, we mimic the principle of hierarchical clustering (Johnson, 1967) in this stage. Concretely, we determine the most similar node pair $(d_{i},d_{j})$ based on their cosine similarity $\text{cos}(\mathbf{e}_{d_{i}},\mathbf{e}_{d_{j}})$, and then update their LCA embedding
@@
This process continues down the similarity list until the remaining highest similarity falls below the threshold $\rho=0.3$. See Appendix J for details of the information flow.
Figure 2: An illustration of hierarchy-guided domain discovery.
Node Scoring. For each node $n\in\mathcal{H}$, we define $S(n)$ to identify which node is a good “candidate” for final selection. Motivated by the information gains (Song and Ying, 2015), $S(n)$ involves three indicators (see in Appendix K), coverage $\mathrm{cov}(n)$, purity $\mathrm{pur}(n)$, and depth $\mathrm{dep}(n)$, via
$S(n)=\exp(\mathrm{pur}(n))+\alpha\cdot\bigl(\mathrm{cov}(n)\times\mathrm{dep}(n)\bigr)$
$=\exp\Bigl(\mathbb{E}_{m\in\mathcal{M}}[\mathrm{sim}(\mathbf{e}_{n},\mathbf{e}_{m})]\Bigr)+\alpha\cdot(\frac{|\mathcal{M}|}{|\mathcal{L}|}\cdot\frac{h}{H}),$ (5)
@@
Case 1. If $S(p)>\max(\{S(c_{i})\}^{r})$, we include parent node $p$ and exclude its children.
Case 2. If $S(p)<\min(\{S(c_{i})\}^{r})$, we discard parent node $p$ and select all children.
@@
Case 3. Otherwise, we tentatively include $p$ but mark it for further resolution in the next step.
Note that, if $p$ is discarded in the above cases, $\{c_{1},\dots,c_{r}\}$ are promoted to become the children of $p$’s parent. After the first iteration on the score matrix, a candidate subset $\mathcal{C}_{0}$ is obtained; however, it remains sub-optimal and requires further evaluation. To this end, we construct a list of tuples $A$ of length $N$, where each element $A[n]:=(p,{c_{1},\dots,c_{r}})$ represents a flagged parent–child pair.
@@
Domain Searching. Given $N$ flagged pairs, the complexity of searching approaches grows exponentially $O(2^{N})$. Inspired by the Beam-Search algorithm (Lowerre, 1976), we get decisions within the top-k pruning candidates $O(kN)$. For each flagged pair, we either (1) unify them by including the parent or (2) retain the children as distinct pruned nodes. Those pairs that have not yet been decided default to retaining parent node to ensure complete coverage. We adopt the Silhouette Score (Shahapure and Nicholas, 2020) as the evaluation metric. Specifically, UdonCare first apply UMAP (McInnes et al., 2018) on leaf node features for dimensionality reduction, and it then calculate the score on this full hierarchical partition, reflecting the global separation quality under the current pruning scheme. This process iterates over all flagged node pairs in order by updating pruning subsets in $\{\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{N}\}$ to make selection with the optimal evaluation result.
Domain Decision. With the pruned selection $\mathcal{C}_{N}$ ($\mathcal{C}^{\prime}$ in Eq. (2)), we update $\mathbf{M}$ by converting patient data into these pruned clusters in $\mathcal{C}_{N}$. If patient $\mathbf{x}^{(i)}$ has at least one leaf code $d$ that descends from the pruned node, we update $\mathbf{M}[i,j]=1$; otherwise, $\mathbf{M}[i,j]=0$. Consequently, each row $\mathbf{M}[i,:]$ represents the domain labels of the $i$-th patient, thereby defining a structured domain partitions over patients.
4.2 Mutual Forward Learning